OpenStudy (minecraftqueen):

Which literal equations are equivalent to k= F/x? Choose all answers that are correct. A. F= kx B. x= kf C. F= k/x D. X= F/k

OpenStudy (minecraftqueen):

ii think its a and b @inowalst is it right

OpenStudy (anonymous):

okay think as a ratios

OpenStudy (anonymous):

ratios means a numbers that are equal to together so k is equal to f/x

OpenStudy (minecraftqueen):

so not right?

OpenStudy (anonymous):

calm down child, there is more to be shown ;)

OpenStudy (anonymous):

so k = f/x now we need to turn the f/x into a simple letter so we times the x on both side k * x = f kx= f

OpenStudy (anonymous):

so It is therefore A

OpenStudy (minecraftqueen):

it says choose all -_-''

OpenStudy (anonymous):

my advice is get rid of the fraction it is always slow people down. ;)

OpenStudy (anonymous):

ahh okay heres a another way k = f/x lets do the f times the f on both side like we did on the x kf = x

OpenStudy (anonymous):

a and b ;)

OpenStudy (minecraftqueen):

CAN U JUST ANSWERE MY FIRST QUESTION IS WHAT I GUEST RIGHT

OpenStudy (minecraftqueen):

thx and stop winking it make u seem like a perv

OpenStudy (anonymous):

No I cant, It is against the Code of Conduct, I am not allowed to tell you the answer right away untill I explained how I got it right.

OpenStudy (anonymous):

No need to be rude. I dont have to report you, do I?

OpenStudy (minecraftqueen):

-_- if u dont want to but idk @mitchal

OpenStudy (minecraftqueen):

i mean idc